OpenStudy (anonymous):

Rationalize the denominator of 4 over 6 minus sqrt 7

jimthompson5910 (jim_thompson5910):

Hint: multiply top and bottom by \(\large 6+\sqrt{7}\) this is the conjugate of the denominator

OpenStudy (anonymous):

\[24\sqrt{7}\] ?

jimthompson5910 (jim_thompson5910):

no but close

jimthompson5910 (jim_thompson5910):

\(\large 4(6+\sqrt{7}) = ??\)

OpenStudy (anonymous):

24+4sqrt7

jimthompson5910 (jim_thompson5910):

better

jimthompson5910 (jim_thompson5910):

\(\large (6-\sqrt{7})(6+\sqrt{7}) = ??\)

OpenStudy (anonymous):

36-sqrt49?

jimthompson5910 (jim_thompson5910):

the square root of 49 is what

OpenStudy (anonymous):

7

jimthompson5910 (jim_thompson5910):

36-sqrt49 turns into 36 - 7 turns into 29

jimthompson5910 (jim_thompson5910):

So the answer is \[\large \frac{4+4\sqrt{7}}{29}\]

OpenStudy (anonymous):

okay, I got it now! Thanks!!!!

jimthompson5910 (jim_thompson5910):

yw

jimthompson5910 (jim_thompson5910):

ooops i meant to say So the answer is \[\large \frac{24+4\sqrt{7}}{29}\] but you got the idea

OpenStudy (anonymous):

can you help me with another?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

3 --- 5-6i

OpenStudy (anonymous):

Rationalize the denominator

jimthompson5910 (jim_thompson5910):

same idea, multiply top and bottom by 5+6i 5+6i is the complex conjugate of 5-6i

OpenStudy (anonymous):

15+18i --------- 25-36i ?

jimthompson5910 (jim_thompson5910):

close, but the denominator should be 25 - 36i^2

OpenStudy (anonymous):

*25-36i^2

jimthompson5910 (jim_thompson5910):

good catch

jimthompson5910 (jim_thompson5910):

i^2 = -1 25 - 36i^2 25 - 36(-1) 25 + 36 61 So 25 - 36i^2 = 61

OpenStudy (anonymous):

so the answer would be 15+18i ---------- 61 ?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

you can break up the fraction if you want, but you're pretty much done

OpenStudy (anonymous):

thanks!!

jimthompson5910 (jim_thompson5910):

np